OpenStudy (anonymous):

What is the solution set of |2x – 3| – 8 = –1?

OpenStudy (anonymous):

you need to understand what the ABSOLUTE VALUE is. The law for ABSOLUTE VALUE is as follows: \[|x| = x\] if \[x \ge 0\] and similarly \[|x| = -x\] if \[x \lt 0\]

OpenStudy (anonymous):

x=5,-2

OpenStudy (anonymous):

dont listen to him, its not right

OpenStudy (anonymous):

so you have to now set up two seperate cases for which x can be POSITIVE _OR_ x can be NEGATIVE. As in the Absolute value law I showed you above.

OpenStudy (anonymous):

following the FIRST law where we are assuming x >= 0 (POSITIVE): solve for 2x-3-8 = -1 x = 5 Plugin a value less than 5 and more than 5 and see what result you get. Construct a number line and plot your signs or a sign table. Then you would find the INTERVAL in which the numbers within the interval suffice the first law whereby x is positive following the SECOND law where we are assuming x < 0 (NEGATIVE): solve for -(2x-3)-8=-1 x = -2 Plugin a value less than -2 and more than and see what result you get. Construct a number line and plot your signs or a sign table. Then you would find the INTERVAL in which the numbers within the interval suffice the first law whereby x is negative. What I suggest you do is rewrite the initial equation by taking the 8 over and doing the same procedure above. Im not going to give you the answer, you do it.